# Instabilities of the Fractionalized Dirac Semimetal in the Kitaev-Kondo Model

### v1

@@ Abstract
We study a honeycomb Kondo lattice model in which Dirac conduction electrons are coupled to a spin-1/2 Kitaev quantum spin liquid. For weak Kondo coupling, the spins fractionalize into Majorana fermions comprising a gapless Dirac mode and three gapped visons. Integrating out the visons to second order in the Kondo coupling yields effective electron–electron interactions, including a local Hubbard repulsion, a spin–spin interaction whose sign depends on the Kitaev exchange, and a vertex coupling electrons to Majorana fermions. We analyze the resulting low-energy field theory using a perturbative renormalization group (RG) scheme, accounting for additional density–density interactions generated under RG. At criticality, electrons decouple from Majorana fermions but all three electron interactions acquire positive values. An analysis of susceptibility exponents reveals that the fractionalized Fermi liquid becomes unstable towards antiferromagnetic order and that superconductivity is disfavored.
@@ I Introduction
In this paper we use a complementary method to investigate the instabilities of the fractionalized half-filled Dirac semimetal of the Kitaev-Kondo model. By integrating out the gapped vison modes we derive an effective low-energy continuum field theory of the gapless Dirac Majorana fermions coupled to Dirac conduction electrons. It contains a Hubbard on-site repulsion and a spin-spin interactions between conduction electrons, as well as a four-fermion vertex that couples Majorana fermions and conduction electrons. We use a perturbative parquet RG analysis [14, 15, 16, 17] to first obtain the scale dependence of the various interaction parameters, and as a second step determine the susceptibility exponents of different order parameters from the RG flow of the conjugate fields at the critical fixed point.
@@ II Model
Although the resulting Hamiltonian is initially quartic in terms of the Majorana fermion operators, it can be solved analytically since the local bond operators $\hat{A}_{ij}^{\gamma}=i\hat{\xi}_{i}^{\gamma}\hat{\xi}_{j}^{\gamma}$ and corresponding plaquette operators, given by the product of bond operators around each hexagon, commute with the Hamiltonian. This results in a free-fermion Hamiltonian with Dirac dispersion $\epsilon_{0}(\mathbf{k})=\pm K|\lambda(\mathbf{k})|$ of the $\hat{\eta}$ Majorana fermions (spinons) and three flat bands $\epsilon_{\gamma}(\mathbf{k})=\pm\Delta$ for the localized $\hat{\xi}^{\gamma}$ Majorana fermions (visons). The energy gap $\Delta$ ($\Delta/|K|\approx 0.525$ [18, 19]) corresponds to the excitation energy of a pair of $Z_{2}$ gauge fluxes [12]. Note that by definition, the Majorana fermions are always at half filling.
@@ III Effective Low-Energy Theory
To derive an effective low-energy field theory, we integrate out the gapped vison modes $\bm{\xi}^{\gamma}$. This gives rise to interactions $\sim J_{K}^{2}$ between the conduction electrons, as well as an interaction between the gapless Majorana fermions and the conduction electrons.
@@
In addition to the interactions $U$, $J$, and $\Gamma$, which arise from integrating out the gapped vison modes, we will also include a density-density interaction $\rho$ between conduction electrons on neighboring sites,
@@ VI Discussion
We have analyzed the instabilities of the fractionalized Fermi liquid that arrises in a Kondo lattice model of Dirac conduction electrons on the half-filled honeycomb lattice coupled to a Kitaev QSL. After fractionalizing the spin-1/2 operators into Majorana fermions, we have integrated out the gapped vison modes to derive an effective low-energy, continuum field theory. The bare theory contains three different interaction vertices, a Hubbard repulsion $U$ and a spin-spin interaction $J$ between Dirac electrons, as well as a coupling $\Gamma$ between conduction electrons and Dirac Majorana fermions. We have analyzed the scale dependence of the interactions, using perturbative parquet RG. In addition to the aforementioned interactions, an electronic density-density repulsion $\rho$ is generated.
@@
Regardless of the ratio of Fermi velocities and the sign of the Kitaev coupling, the RG flow is always controlled by the same critical fixed point, at which the Dirac electrons decouple from the Majorana fermions and all three electron-electron interaction have finite positive values. To determine the type of symmetry breaking we have derived the RG equations of the fields conjugate to CDW, SDW and SC order, from which we determined the field scaling exponents at the critical point and the susceptibility exponents. For the SC order parameter we have focused on singlet pairing across neighboring sites with a spatial $d+id$ form factor [13]. While the antiferromagnetic spin-spin interaction $J$ stabilizes antiferromagnetism and SC in almost equal terms, the AFM ordering is strongly enhanced by the Hubbard repulsion $U$ whereas SC is destabilized by the repulsive density-density interaction $\rho$. Our results show that the only divergent susceptibility is toward antiferromagnetic (SDW) order.
@@
While the perturbative parquet RG is the method of choice to identify the leading symmetry-breaking instability in the presence of multiple competing interactions [14, 15, 16, 17], the critical exponents for the correlation length $\nu=1$ and SDW susceptibility $\gamma=2/3$ obtained this way are only crude estimates. Ultimately, the SDW transition at half filling should falls into the Gross-Neveu-Heisenberg universality class in 2+1 dimensions. By combining results from Padé approximations of $4-\epsilon$ expansions [23], $1/N$ expansions in the number of fermion flavors [24], and functional RG [25], estimates of $1/\nu\approx 0.886$ for the inverse correlation length exponent and $\eta_{\phi}\approx 1.035$ for anomalous dimension of the order parameter field were obtained [26]. Assuming hyperscaling, this would imply $\gamma=(2-\eta_{\phi})\nu\approx 1.089$.
@@ Acknowledgements.
J.L. is grateful for discussions with Cristian Batista, Claudio Castelnovo, and Cecilie Glittum. F.K. acknowledges fruitful discussions with Chris Hooley.

### v2

@@ Abstract
We study a honeycomb Kondo lattice model in which Dirac conduction electrons are coupled to a spin-1/2 Kitaev quantum spin liquid. For weak Kondo coupling, the spins fractionalize into Majorana fermions comprising a gapless Dirac mode and three gapped visons. In second order perturbation theory, the Kondo coupling gives rise to local Hubbard repulsions and spin–spin interactions between conduction electrons, as well as a vertex coupling electrons to gapless Majorana fermions. We analyze the resulting low-energy field theory using a perturbative renormalization group (RG) scheme, accounting for additional density–density interactions generated under RG. At criticality, electrons decouple from Majorana fermions but all three electron interactions acquire positive values. An analysis of susceptibility exponents reveals that the fractionalized Fermi liquid becomes unstable towards antiferromagnetic order and that superconductivity is disfavored.
@@ I Introduction
In this paper we use a complementary method to investigate the instabilities of the fractionalized half-filled Dirac semimetal of the Kitaev-Kondo model. Treating the Kondo coupling in second-order perturbation theory, we derive an effective low-energy continuum field theory of the gapless Dirac Majorana fermions coupled to Dirac conduction electrons. It contains a Hubbard on-site repulsion and a spin-spin interactions between conduction electrons, as well as a four-fermion vertex that couples Majorana fermions and conduction electrons. We use a perturbative parquet RG analysis [14, 15, 16, 17] to first obtain the scale dependence of the various interaction parameters, and as a second step determine the susceptibility exponents of different order parameters from the RG flow of the conjugate fields at the critical fixed point.
@@ II Model
Although the resulting Hamiltonian is initially quartic in terms of the Majorana fermion operators, it can be solved analytically since the local bond operators $\hat{A}_{ij}^{\gamma}=i\hat{\xi}_{i}^{\gamma}\hat{\xi}_{j}^{\gamma}$ and corresponding plaquette operators, given by the product of bond operators around each hexagon, commute with the Hamiltonian. This results in a free-fermion Hamiltonian with Dirac dispersion $\epsilon_{0}(\mathbf{k})=\pm K|\lambda(\mathbf{k})|$ of the $\hat{\eta}$ Majorana fermions (spinons) and three flat bands $\epsilon_{\gamma}(\mathbf{k})=\pm\Delta$ for the localized $\hat{\xi}^{\gamma}$ Majorana fermions (visons) with an energy gap $\Delta/|K|\approx 0.525$ [18, 19, 13]. Note that by definition, the Majorana fermions are always at half filling.
@@ III Effective Low-Energy Theory
To derive an effective low-energy field theory, we use second order perturbation theory to obtain the effective interactions $\sim J_{K}^{2}$ between conduction electrons, as well as an interaction between the gapless Majorana fermions and the conduction electrons.
@@
In general, the Kitaev exchange between local moment spins is mediated through a super-exchange mechanism and therefore expected to be weak compared to typical electronic energy scales. For example, ferromagnetic Kitaev exchanges in the range 5-20 meV have been estimated for $\alpha$-RuCl3 [20, 21] and Na2IrO3 [22], while certain Co-based honeycomb materials were found to exhibit an antiferromagnetic Kitaev exchange of 2-3 meV [23]. For comparison, the nearest-neighbor tight-binding parameter for graphene is $t\approx 2.7$ eV [24]. In real materials one should therefore expect a Fermi-velocity ratio $c/v$ of the order of $10^{-3}-10^{-2}$, implying that $\Gamma\ll U$. However, we will theoretically explore the full range of $c/v$ and demonstrate that even in the regime of very strong coupling $\Gamma$ between Majorana fermions and conduction electrons the critical behaviour will remain the same.
In addition to the interactions $U$, $J$, and $\Gamma$ we will also include a density-density interaction $\rho$ between conduction electrons on neighboring sites,
@@ IV Renormalization-Group Analysis
The four interactions $U$, $J$, $\rho$ and $\Gamma$ form a basis set for the subspace of interactions explored under the RG since it is not possible to express one of the interactions as a linear combination of the other three. However, by using a Fierz identity one may express the density-density interaction $\rho$ in terms of the spin-spin interaction $J$ and an additional pair-hopping vertex. Such ambiguities could lead to biased results when carrying out a Hubbard-Stratonovich or mean-field decoupling but wouldn’t affect our perturbative RG analysis.
@@ V Field renormalization and susceptibility exponents
The Landau free-energy analysis of different SC order parameters of Ref. [13] is fairly general and should also apply to electron fillings close to the Dirac point. One would therefore expect that the chiral $d\pm id$ state remains the supeconducting state with the lowest free energy at half filling. However, for completeness, we have checked superconducting states with other pairing symmetries. As one might expect, we found that triplet pairing is disfavoured by antiferromagnetic $J$, while on-site singlet pairing is strongly suppressed by the Hubbard repulsion $U$.
@@ VI Discussion
We have analyzed the instabilities of the fractionalized Fermi liquid that arrises in a Kondo lattice model of Dirac conduction electrons on the half-filled honeycomb lattice coupled to a Kitaev QSL. After fractionalizing the spin-1/2 operators into Majorana fermions, we have used second-order perturbation theory to obtain the interactions between the gapless fermion modes in the effective low-energy, continuum field theory. The bare theory contains three different interaction vertices, a Hubbard repulsion $U$ and a spin-spin interaction $J$ between Dirac electrons, as well as a coupling $\Gamma$ between conduction electrons and Dirac Majorana fermions. We have analyzed the scale dependence of the interactions, using perturbative parquet RG. In addition to the aforementioned interactions, an electronic density-density repulsion $\rho$ is generated.
Regardless of the ratio of Fermi velocities and the sign of the Kitaev coupling, the RG flow is always controlled by the same critical fixed point, at which the Dirac electrons decouple from the Majorana fermions and all three electron-electron interaction have finite positive values. This results suggests that it is justified to integrate out the Majorana fermions entirely, as done in Ref. [13]. However, we stress that the velocity ratio $c/v$ could potentially renormalize beyond one-loop order, resulting in a different flow of the coupling $\Gamma$ between conduction electrons and Dirac Majorana fermions. Velocity renormalization is a well known phenomenon of strongly coupled GNY theories and perturbative RG, even beyond one-loop order, cannot unambiguously rule out the existence of a novel GNY critical fixed point with finite coupling to Majorana fermions.
@@
To determine the type of symmetry breaking we have derived the RG equations of the fields conjugate to CDW, SDW and SC order, from which we determined the field scaling exponents at the critical point and the susceptibility exponents. For the SC order parameter we have focused on singlet pairing across neighboring sites with a spatial $d+id$ form factor [13]. While the antiferromagnetic spin-spin interaction $J$ stabilizes antiferromagnetism and SC in almost equal terms, the AFM ordering is strongly enhanced by the Hubbard repulsion $U$ whereas SC is destabilized by the repulsive density-density interaction $\rho$. Our results show that the only divergent susceptibility is toward antiferromagnetic (SDW) order.
@@
While the perturbative parquet RG is the method of choice to identify the leading symmetry-breaking instability in the presence of multiple competing interactions [14, 15, 16, 17], the critical exponents for the correlation length $\nu=1$ and SDW susceptibility $\gamma=2/3$ obtained this way are only crude estimates. Ultimately, the SDW transition at half filling should falls into the Gross-Neveu-Heisenberg universality class in 2+1 dimensions. From interpolation between series expansions near lower and upper critical dimensions [28] with previous results from fourth-order $4-\epsilon$ [29] and second-order $1/N$ [30] expansions, functional RG [31, 32], as well as quantum Monte Carlo simulations [33, 34, 35, 36, 37, 38, 39, 40, 41, 42] estimates of $1/\nu\approx 0.83$ for the inverse correlation length exponent and $\eta_{\phi}\approx 1.01$ for the anomalous dimension of the order parameter field were obtained for $N=4$ (valley and spin) 2-component Dirac fermion fields [28]. Assuming hyperscaling, this would imply $\gamma=(2-\eta_{\phi})\nu\approx 1.19$.
@@
To better understand the instabilities of different types of fractionalized Fermi liquids towards exotic superconductivity and other broken symmetry states it will be important to investigate other toy models that host fractionalized FL phases. In a very recent study [45] it was demonstrated that when the so called “Yao-Lee Model” [46], a spin-orbital generalisation of the Kitaev model, is Kondo-coupled to conduction electrons, the exchange of Majorana spinons can drive a (fractionalized) topological $d+id$ pairing instability in the weak-coupling regime. In future studies it would be interesting to allow for Kondo coupling not only to the SU(2) symmetric local moment spins but also to the Kitaev orbital pseudo-spin sector and to systematically investigate the dependence on the electron filling.
A frustrated Kondo lattice model with an FL phase that is s amenable to sign free auxiliary-field quantum Monte Carlo simulations was constructed [5] by coupling Dirac electrons on the honeycomb lattice to spin-1/2 degrees of freedom on the Kagome lattice. The interactions between the spins are chosen along the lines of the Balents-Fisher-Girvin model [47] which is known to host a $Z_{2}$ spin-liquid and a ferromagnetic phase. While the authors characterized the FL phase in great detail through spectral functions and mutual information between electrons and spins, they did not investigate possible superconducting instabilities.
Studies of various toy models have shown that fractionalized Fermi liquids naturally arise when quantum spin liquids are weakly coupled to conduction electrons in a Kondo lattice setting. The emergent fermions in such FL phases mediate effective electron-electron interactions that can drive instabilities towards exotic superconductivity or other symmetry-broken phases. The emergent fermions therefore act as pairing glue similar to the phonons in conventional BCS superconductivity. Our work shows that at the quantum critical point of the FL phase of the half-filled Kitaev-Kondo model the emergent Majorana fermions decouple from the conduction electrons. It remains an intriguing question wether this behavior is generic or if certain models can support quantum critical points at which emergent fermions and conduction electrons remain strongly coupled, opening up the possibility for the formation of even more exotic quantum phases.
@@ Acknowledgements.
J.L. is grateful for discussions with Cristian Batista, Claudio Castelnovo, and Cecilie Glittum. F.K. acknowledges fruitful discussions with Chris Hooley, Lukas Janssen, and Hong Yao.